OpenStudy (anonymous):

1. -3(6a-2)=114 2. 5+6(1-4x)=-109

OpenStudy (anonymous):

-3(6a-2)=114 u distribute -3 to 6a and to -2 and multiply to get -18a+6=114 then sub tract 6 from both sides -18a+6=114 -6 -6 ------------ 108 then divid -18 from both sides -18a = 108 --- --- -18 -18 u get -6 so a equals -6 (: hope i helped

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

ur welcome (: